### v1

# Phantom crossing or dark interaction?

Abstract
Recent results from DESI BAO measurements, together with Planck CMB and Pantheon+ data, suggest that there may be a ‘phantom’ phase ($w_{\rm de}<-1$) in the expansion of the Universe. This inference follows when the $w_{0},w_{a}$ parametrization for the dark energy equation of state $w_{\rm de}$ is used to fit the data. Since phantom dark energy in general relativity is unphysical, we investigate the possibility that the phantom behaviour is not intrinsic, but effective – due to a non-gravitational interaction between dark matter and non-phantom dark energy. To this end, we assume a physically motivated thawing quintessence-like form of the intrinsic dark energy equation of state $w_{\rm de}$. Then we use a $w_{0},w_{a}$ model for the effective equation of state of dark energy. We find that the data favours a phantom crossing for the effective dark energy, but only at low significance. The intrinsic equation of state of dark energy is non-phantom, without imposing any non-phantom priors. A nonzero interaction is favoured at more than $3\sigma$ at $z\sim 0.3$. The energy flows from dark matter to dark energy at early times and reverses at later times.
July 24, 2025
1 Introduction
The origin and dynamics of dark matter and dark energy, which together account for approximately 95% of the total current energy content of the Universe, remain among the greatest unsolved questions in modern cosmology. In the standard $\Lambda$CDM model, dark energy is modelled as a cosmological constant $\Lambda$, and dark matter is treated as a cold, pressureless component, with no non-gravitational interaction between them [1, 2, 3].
While this framework provides an excellent fit to a wide range of cosmological observations, it faces several theoretical and empirical challenges, including tensions in measurements of the Hubble constant and large-scale structure growth, the cosmic coincidence problem, and the cosmological constant problem (see e.g. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]). These issues motivate the exploration of alternative models, including those in which dark matter and dark energy interact non-gravitationally (see e.g. [14, 15, 16, 17, 18, 19, 20, 21]).
These models introduce an interaction term $Q(z)$ in the background energy conservation equations, allowing for energy exchange between the dark components. The sign of $Q$ determines the direction of energy flow: $Q>0$ corresponds to energy transfer from dark energy to dark matter, and vice-versa for $Q<0$. Several studies have shown that dark interaction models can address some of the potential issues in cosmology, such as the Hubble tension and the crossing of the phantom divide [22, 23, 24, 25, 26].
However, a major challenge in analysing dark interaction models arises from the degeneracy between the interaction $Q(z)$ and the dark energy equation of state parameter $w_{\rm de}(z)$. Both quantities influence the expansion history of the Universe, making it impossible to study their effects individually, using background observations alone, without additional assumptions [22, 27, 28, 26, 29, 30, 31, 32]. Model-dependent approaches assume specific functional forms for the interaction, but typically these are not physically well-motivated (see e.g. [15]). Here we do not make assumptions about $Q(z)$ – instead, we adopt a physically motivated form for the intrinsic $w_{\rm de}(z)$ and the $w_{0}w_{a}$ parametrization for the effective equation of state of dark energy $w_{\rm de}^{\rm eff}(z)$, to break the degeneracy between $w_{\rm de}(z)$ and $Q(z)$ (see e.g. [27, 29, 30] for related work).
Recent advances in observational cosmology, especially from the Dark Energy Spectroscopic Instrument Data Release 2 (DESI DR2) baryon acoustic oscillation (BAO) measurements [23, 24], the Planck cosmic microwave background (CMB) data [33], and the Pantheon+ supernova compilation [34], provide high precision for probing the dark sector and we study the dark interaction in light of these observations.
The paper is organised as follows. In section 2, we present the theoretical framework for the interacting dark sector and derive the key equations governing the interaction. The observational data are briefly reviewed in section 3 and the results are presented in section 4. We conclude in section 5 with a discussion of the implications of our findings.
2 Background dynamics of dark interaction
We consider a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) universe governed by general relativity, where the dark sector consists of dark matter and dark energy with equation of state $w_{\rm de}=p_{\rm de}/\rho_{\rm de}$. The total energy density satisfies the Friedmann equation:
$H(z)^{2}=\frac{8\pi G}{3}\left[\rho_{\rm dm}(z)+\rho_{\rm de}(z)+\rho_{\rm b}(z)+\rho_{\rm r}(z)\right]\,,$ (2.1)
where $\rho_{\rm a}(z)$ denotes the energy densities of dark matter, dark energy, baryons, and radiation, respectively, and $H(z)$ is the Hubble parameter.
In the presence of a non-gravitational interaction between dark matter and dark energy, the standard conservation equations are modified:
$\dot{\rho}_{\rm dm}+3H\rho_{\rm dm}$ $=Q\,,$ (2.2)
$\dot{\rho}_{\rm de}+3H\big{(}1+w_{\rm de}\big{)}\rho_{\rm de}$ $=-Q\,.$ (2.3)
Radiation and baryons are assumed to be separately conserved at late times:
$\dot{\rho}_{\rm b}+3H\rho_{\rm b}=0,\quad\dot{\rho}_{\rm r}+4H\rho_{\rm r}=0\,.$ (2.4)
Equation 2.2 and Equation 2.3 can be equivalently described by a non-interacting model with effective equations of state:
$w^{\rm eff}_{\rm dm}(z)$ $=-\frac{Q(z)}{3H(z)\rho_{\rm dm}(z)}\,,$ (2.5)
$w^{\rm eff}_{\rm de}(z)$ $=w_{\rm de}(z)+\frac{Q(z)}{3H(z)\rho_{\rm de}(z)}\,,$ (2.6)
corresponding to
$\dot{\rho}_{\rm dm}+3H\big{(}1+w^{\mathrm{eff}}_{\rm dm}\big{)}\rho_{\rm dm}$ $=0\,,$ (2.7)
$\dot{\rho}_{\rm de}+3H\big{(}1+w^{\mathrm{eff}}_{\rm de}\big{)}\rho_{\rm de}$ $=0\,.$ (2.8)
Equation 2.5 and Equation 2.6 show that among the four quantities $w_{\rm de}$, $w_{\rm de}^{\rm eff}$, $Q$, and $w_{\rm dm}^{\rm eff}$, only two are independent. Hence the interaction model can be fully specified by any two of these quantities. Here we specify the model through $w_{\rm de}$ and $w_{\rm de}^{\rm eff}$, which fixes $Q$ and $w_{\rm dm}^{\rm eff}$.
Equation 2.1 is rewritten as
$H(z)=H_{0}\left[\Omega_{\rm r,0}(1+z)^{4}+\Omega_{\rm b,0}(1+z)^{3}+\Omega_{{\rm dm},0}f_{\rm dm}(z)+\Omega_{{\rm de},0}f_{\rm de}(z)\right]^{1/2},$ (2.9)
where $\Omega_{\rm a,0}=\rho_{\rm a,0}/\rho_{\rm crit,0}$ and $\rho_{\rm crit,0}=3H_{0}^{2}/(8\pi G)$. The normalized energy densities of dark matter and dark energy are $f_{\rm dm}(z)=\rho_{\rm dm}(z)/\rho_{\rm dm,0}$ and $f_{\rm de}(z)=\rho_{\rm de}(z)/\rho_{\rm de,0}$ and $\Omega_{\rm r,0}+\Omega_{\rm b,0}+\Omega_{{\rm dm},0}+\Omega_{{\rm de},0}=1$. In the standard non-interacting case $Q=0$, we have $f_{\rm dm}=(1+z)^{3}$, $f_{\rm de}=\exp{\left[3\,\int_{0}^{z}\,\mathrm{d}z^{\prime}[1+w_{de}(z^{\prime})]/(1+z^{\prime})\right]}$. The present-day radiation density parameter is
$\Omega_{\rm r,0}=\frac{\Omega_{\rm m,0}}{1+z_{\rm eq}},$ (2.10)
where the matter-radiation equality redshift in standard early-time physics is well approximated as [35]
$z_{\rm eq}\approx 2.5\times 10^{4}\,\Omega_{\rm m,0}h^{2}\left(\frac{T_{\rm CMB}}{2.7\,\mathrm{K}}\right)^{-4}.$ (2.11)
Here $h=H_{0}/(100\,\mathrm{km\,s^{-1}Mpc^{-1}})$ and $T_{\rm CMB}\approx 2.7255$ K is the background CMB temperature.
The degeneracy between $Q(z)$ and $w_{\rm de}(z)$ means that it is not possible, from background data, to reconstruct either of these independently without assuming any model or parametrization. We aim to reconstruct $Q(z)$ by using a physically motivated parametrization of $w_{\rm de}(z)$, corresponding to thawing quintessence models. In these models, a minimally coupled scalar field is frozen at early times due to Hubble friction, effectively behaving as a cosmological constant with $w_{\rm de}\approx-1$. It evolves later, leading to a deviation from a cosmological constant at late times with increasing $w_{\rm de}>-1$. We adopt the analytic from [23]:
$w_{\rm de}(z)=-1+(1+w_{\rm q})(1+z)^{-1}\left[\frac{3}{2+(1+z)^{3}}\right]^{2/3}\quad\mbox{where}\quad w_{\rm q}=w_{\rm de}(0)\,.$ (2.12)
Note that the above form corresponds to thawing quintessence only when $w_{\rm q}>-1$. On the other hand, it behaves like phantom dark energy for $w_{\rm q}<-1$. This is the intrinsic dark energy equation state. However, observations will measure the effective equation of state, for which we assume the standard $w_{0},w_{a}$ parametrization
$w^{\rm eff}_{\rm de}(z)=w_{0}+w_{a}\frac{z}{1+z}\,.$ (2.13)
Then $Q$ is determined by $w_{\rm de}$ and $w_{\rm de}^{\rm eff}$. Our aim now is as follows:
  • to show that the data allows for an interaction with intrinsic and effective equations of state, of the forms Equation 2.12 and Equation 2.13;
  • to then find the level of interaction required for this to be the case.
From Equation 2.8, we have
$f_{\rm de}(z)=\exp\left\{3\!\int_{0}^{z}\left[1+w^{\mathrm{eff}}_{\rm de}(z^{\prime})\right]\frac{\mathrm{d}z^{\prime}}{1+z^{\prime}}\right\}=(1+z)^{3(1+w_{0}+w_{a})}\,\exp{\left(\!-3w_{a}\,\frac{z}{1+z}\right)}\,.$ (2.14)
Then Equation 2.6 gives
$Q(z)=3H(z)\rho_{\rm crit,0}\,\Omega_{\rm de,0}\big{[}w^{\rm eff}_{\rm de}(z)-w_{\rm de}(z)\big{]}f_{\rm de}(z)\,,$ (2.15)
and it follows from Equation 2.5 that
$w^{\rm eff}_{\rm dm}(z)=\big{[}w_{\rm de}(z)-w^{\rm eff}_{\rm de}(z)\big{]}\,\frac{\Omega_{\rm de,0}\,f_{\rm de}(z)}{\Omega_{\rm dm,0}\,f_{\rm dm}(z)}\,.$ (2.16)
Substituting into Equation 2.7, we find a differential equation for $f_{\rm dm}$:
$\dfrac{{\rm d}f_{\rm dm}}{{\rm d}z}=\frac{3}{1+z}\left[f_{\rm dm}+\big{(}w_{\rm de}-w^{\rm eff}_{\rm de}\big{)}\,\frac{\Omega_{\rm de,0}\,f_{\rm de}}{\Omega_{\rm dm,0}}\right]\,,$ (2.17)
with solution
$f_{\rm dm}(z)$ $=(1+z)^{3}\Bigg{\{}1+3\frac{\Omega_{\rm de,0}}{\Omega_{\rm dm,0}}\,\int_{1}^{(1+z)^{-1}}\,\mathrm{d}a\bigg{[}\left(1+w_{0}+w_{a}\right)-w_{a}a$
$\qquad\qquad-(1+w_{\rm q})\,a^{2}\left(\frac{3}{2a^{3}+1}\right)^{2/3}\bigg{]}\,{a}^{-(1+3w_{0}+3w_{a})}\,\exp{\big{[}-3w_{a}(1-a)\big{]}}\Bigg{\}}\,.$ (2.18)
We can then infer the Hubble parameter from Equation 2.9 and constrain the parameters $w_{\rm q}$, $w_{0}$, $w_{a}$ and other cosmological parameters.
3 Observational Data
In order to constrain parameters and reconstruct the dark sector interaction function $Q(z)$, we use three complementary cosmological datasets:
  • DESI DR2 BAO [24]: Measurements of BAO from DESI DR2 provide high-precision determinations of the comoving angular diameter distance $D_{M}(z)$, the Hubble distance $D_{H}(z)=c/H(z)$, and the volume-averaged distance $D_{V}(z)$. These are reported in units of the comoving sound horizon at the drag epoch $r_{d}$ (i.e. uncalibrated):
    $\widetilde{D}_{M}(z)=\frac{D_{M}(z)}{r_{d}},\qquad\widetilde{D}_{H}(z)=\frac{D_{H}(z)}{r_{d}},\qquad\widetilde{D}_{V}(z)=\frac{D_{V}(z)}{r_{d}},$ (3.1)
    where
    $D_{V}(z)=\left[z\,D_{M}(z)^{2}\,D_{H}(z)\right]^{1/3},\qquad D_{M}(z)=\int_{0}^{z}\frac{c}{H(z^{\prime})}\,\mathrm{d}z^{\prime}\,.$ (3.2)
  • Planck CMB [24]: We use compressed likelihoods from Planck 2018 measurements of three key early-Universe parameters: the baryon density $\omega_{\rm b}=\Omega_{\rm b,0}h^{2}$, the matter density $\omega_{m}=\Omega_{m,0}h^{2}$ and the angular size of the sound horizon at recombination $\theta_{*}=r_{*}/D_{M}(z_{*})$, where $r_{*}$ is the comoving sound horizon at photon decoupling and $z_{*}\approx 1089.8$. We adopt a Gaussian prior on $(\theta_{*},\omega_{\rm b},\omega_{m})$ with covariance matrix ${\bm{C}}$ and mean values $\mu$ from [24]:
    $\mu(\theta_{\star},\omega_{b},\omega_{m})=\left(\begin{array}[]{c}0.01041\\ 0.02223\\ 0.14208\\ \end{array}\right)\!,~{}~{}\bm{C}=10^{-9}\,\left(\begin{array}[]{ccc}0.006621{}&~{}0.12444{}&~{}-1.1929\\ 0.12444{}&~{}21.344{}&~{}-94.001\\ -1.1929{}&~{}-94.001{}&~{}1488.4\\ \end{array}\right)\!.$ (3.9)
    Note that the comoving sound horizon at redshift $z$ is given by:
    $r_{s}(z)=\int_{z}^{\infty}\frac{c_{s}(z^{\prime})}{H(z^{\prime})}\,\mathrm{d}z^{\prime}\,,$ (3.10)
    where the sound speed in the baryon-photon fluid is given by [35]:
    $\frac{c_{s}(z)^{2}}{c^{2}}=\frac{1}{3}\left[1+\frac{3\rho_{\rm b}(z)}{4\rho_{\gamma}(z)}\right]^{-1}=\frac{1}{3}\left[1+\frac{3}{4\Omega_{\gamma,0}\,h^{2}}\,\frac{\Omega_{\rm b,0}h^{2}}{1+z}\right]^{-1}\,.$ (3.11)
    We compute the sound horizon at the drag epoch, $r_{d}=r_{s}(z_{d})$, where $z_{d}\sim 1060$, by integrating numerically. The CMB temperature fixes the photon density today via:
    $\frac{3}{4\Omega_{\gamma,0}h^{2}}\approx 31500\left(\frac{T_{\text{CMB}}}{2.7\,\text{K}}\right)^{-4}\,.$ (3.12)
  • Pantheon+ SNIa [34]: We include type Ia supernova data from the Pantheon+ sample, which provides measurements of the apparent peak magnitude $m_{B}(z)$. These are related to the luminosity distance $D_{L}(z)$ via:
    $m_{B}(z)=M_{B}+5\log_{10}\left[\frac{D_{L}(z)}{\text{Mpc}}\right]+25,\qquad D_{L}(z)=(1+z)D_{M}(z),$ (3.13)
    where $M_{B}$ is the absolute peak magnitude of the type Ia supernovae.
4 Results and Discussion
Figure 1: Triangle plot for DESI DR2 BAO + CMB + Pantheon+ constraints.
Using the data combination DESI DR2 BAO + CMB + Pantheon+, we find the contour plots in Figure 1, corresponding to the posterior distributions of the parameters. The mean and standard deviation values are given in the plot. These values are used to propagate uncertainties in the quantities of interest with all possible cross-correlations, which we have not explicitly quoted.
Figure 2: Intrinsic (left) and effective (right) equations of state for dark energy, with uncertainties. The $\Lambda$ model is the straight red line.
The intrinsic and effective equations of state of the dark energy $w_{\rm de}(z)$ and $w_{\rm de}^{\rm eff}(z)$, are shown in Figure 2. Note that for $w_{\rm de}$ we did not impose a $w_{\rm q}\geq-1$ prior. Although the parametrization Equation 2.12 does not include this condition (unlike the actual quintessence model with scalar field evolution), the data shows a preference for non-phantom behaviour of the intrinsic dark energy (see [28, 36] for a similar result in the non-interacting case). This means that thawing quintessence with interaction term is a good fit to the current data.
[…]
Figure 3: Evolution of the interaction, relative to the contribution of intrinsic dark energy.
[…]
Figure 4: Evolution of the effective dark matter equation of state (left) and the dark sector energy densities (right).
For dark matter, an interaction can be equivalently described by an effective non-interacting and non-cold dark matter, i.e., with effective nonzero equation of state, $w_{\rm dm}^{\rm eff}(z)$ [37]. This is shown in the left panel of Figure 4. There is no significant evidence for nonzero $w_{\rm dm}^{\rm eff}(z)$, except for $\sim 2\sigma$ evidence for $w_{\rm dm}^{\rm eff}<0$ at $z\sim 0.3$.
Finally, in the right panel of Figure 4, we compare $\Omega_{\rm dm}$ and $\Omega_{\rm de}$ to see that the dark matter and dark energy equality happens at $z\sim 0.5$. All the significant changes happen close to this redshift, as seen in the previous plots. This might be a hint of an intrinsic correlation between dark matter-dark energy equality and the redshift where interaction energy transfer changes sign.
5 Conclusion
The $w_{0},w_{a}$ parametrization of the dark energy equation state leads to claims of phantom crossing on the basis of the latest cosmological data. We tested the idea that phantom behaviour – which is physically inconsistent in general relativity – is not real, but only apparent. In other words, it could be a signal of interaction in the dark sector. For this purpose, we assumed a physically self-consistent model of intrinsic dark energy, in the form of thawing quintessence-like parametrization, and we assumed a $w_{0},w_{a}$ parametrization of the effective equation of state of dark energy. Then we reconstructed the redshift dependence of the interaction function, $Q(z)$.
The parametrization of the intrinsic equation of state of dark energy behaves like thawing quintessence for $w_{\rm q}>-1$. On the other hand, it has phantom behavior for $w_{\rm q}<-1$. We did not impose a $w_{\rm q}>-1$ prior, but we still found non-phantom values of $w_{\rm de}(z)$ from current data, hinting that thawing quintessence models may be a good fit to the data in dark interaction models. The effective equation of state of dark energy suggests that there is no significant evidence for phantom crossing even in $w_{\rm de}^{\rm eff}$, which is assumed to be a $w_{0},w_{a}$ parametrization.
We find that the interaction function $Q(z)$ shows a preference for negative values (although not significant) at higher redshifts ($z\gtrsim 0.5$) and changes sign around $z\sim 0.5$ to become positive at lower redshifts. This suggests an energy flow from dark matter to dark energy in the early Universe and energy flow from dark energy to dark matter in the late Universe. Interestingly, this transition happens close to the era of dark matter-dark energy equality.
[…]
In summary, we find interacting dark energy models can be viable alternatives to both the standard $\Lambda$CDM model and phantom dark energy scenarios, while the intrinsic dark energy equation of state is non-phantom.
Acknowledgments
The authors are supported by the South African Radio Astronomy Observatory and the National Research Foundation (grant no. 75415).

### v2

@@
 The energy flows from dark matter to dark energy at early times and reverses at later times.
-July 24, 2025
 1 Introduction
-The origin and dynamics of dark matter and dark energy, which together account for approximately 95% of the total current energy content of the Universe, remain among the greatest unsolved questions in modern cosmology.
-In the standard $\Lambda$CDM model, dark energy is modelled as a cosmological constant $\Lambda$, and dark matter is treated as a cold, pressureless component, with no non-gravitational interaction between them [1, 2, 3].
+[…]
 While this framework provides an excellent fit to a wide range of cosmological observations, it faces several theoretical and empirical challenges, including tensions in measurements of the Hubble constant and large-scale structure growth, the cosmic coincidence problem, and the cosmological constant problem (see e.g.
-[4, 5, 6, 7, 8, 9, 10, 11, 12, 13]).
+[64, 57, 30, 55, 1, 20, 21, 61, 25, 12, 42, 46, 47, 48, 51, 17, 4, 40, 8]).
 These issues motivate the exploration of alternative models, including those in which dark matter and dark energy interact non-gravitationally (see e.g.
-[14, 15, 16, 17, 18, 19, 20, 21]).
+[63, 44, 56, 41, 32, 35, 34, 22, 54, 18, 58, 29, 19, 14, 50, 9]).
 These models introduce an interaction term $Q(z)$ in the background energy conservation equations, allowing for energy exchange between the dark components.
@@
 The total energy density satisfies the Friedmann equation:
-$H(z)^{2}=\frac{8\pi G}{3}\left[\rho_{\rm dm}(z)+\rho_{\rm de}(z)+\rho_{\rm b}(z)+\rho_{\rm r}(z)\right]\,,$ (2.1)
+$H(z)^{2}=\frac{8\pi G}{3}\big[\rho_{\rm dm}(z)+\rho_{\rm de}(z)+\rho_{\rm b}(z)+\rho_{\rm r}(z)\big]\,$, (2.1)
 where $\rho_{\rm a}(z)$ denotes the energy densities of dark matter, dark energy, baryons, and radiation, respectively, and $H(z)$ is the Hubble parameter.
@@
 We aim to reconstruct $Q(z)$ by using a physically motivated parametrization of $w_{\rm de}(z)$, corresponding to thawing quintessence models.
-In these models, a minimally coupled scalar field is frozen at early times due to Hubble friction, effectively behaving as a cosmological constant with $w_{\rm de}\approx-1$.
+In these models, a minimally coupled scalar field is frozen at early times due to Hubble friction, effectively behaving as a cosmological constant with $w_{\rm de}\approx-1$ [28].
 It evolves later, leading to a deviation from a cosmological constant at late times with increasing $w_{\rm de}>-1$.
@@
 with solution
-$f_{\rm dm}(z)$ $=(1+z)^{3}\Bigg{\{}1+3\frac{\Omega_{\rm de,0}}{\Omega_{\rm dm,0}}\,\int_{1}^{(1+z)^{-1}}\,\mathrm{d}a\bigg{[}\left(1+w_{0}+w_{a}\right)-w_{a}a$
-$\qquad\qquad-(1+w_{\rm q})\,a^{2}\left(\frac{3}{2a^{3}+1}\right)^{2/3}\bigg{]}\,{a}^{-(1+3w_{0}+3w_{a})}\,\exp{\big{[}-3w_{a}(1-a)\big{]}}\Bigg{\}}\,.$ (2.18)
+$f_{\rm dm}(z)$ $=(1+z)^{3}\Bigg\{1+3\frac{\Omega_{\rm de,0}}{\Omega_{\rm dm,0}}\,\int_{1}^{(1+z)^{-1}}\,\mathrm{d}a\bigg[\left(1+w_{0}+w_{a}\right)-w_{a}a$ (2.18)
+$\qquad\qquad\quad~-(1+w_{\rm q})\,a^{2}\left(\frac{3}{2a^{3}+1}\right)^{2/3}\bigg]\,{a}^{-(1+3w_{0}+3w_{a})}\,\exp{\big[-3w_{a}(1-a)\big]}\Bigg\}\,.$
 We can then infer the Hubble parameter from Equation 2.9 and constrain the parameters $w_{\rm q}$, $w_{0}$, $w_{a}$ and other cosmological parameters.
+The effective equation of state of all the components of the model can be approximated as
+$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{w_{\rm tot}^{\rm eff}(z)=E^{-2}(z)\left[w^{\rm eff}_{\rm de}(z)\,\Omega_{{\rm de},0}\,f_{\rm de}(z)+w^{\rm eff}_{\rm dm}(z)\,\Omega_{{\rm dm},0}\,f_{\rm dm}(z)+\frac{1}{3}\Omega_{\rm r,0}(1+z)^{4}\right]\,,}$} (2.19)
+where $E(z)=H(z)/H_{0}$. At late times, we can neglect the radiation components and hence the last term in this equation.
 3 Observational data
@@
 Figure 4: Evolution of the effective dark matter equation of state (left) and the dark sector energy densities (right).
-For dark matter, an interaction can be equivalently described by an effective non-interacting and non-cold dark matter, i.e., with effective nonzero equation of state, $w_{\rm dm}^{\rm eff}(z)$ [37].
+For dark matter, an interaction can be equivalently described by an effective non-interacting and non-cold dark matter, i.e., with effective nonzero equation of state, $w_{\rm dm}^{\rm eff}(z)$ [11, 33].
 This is shown in the left panel of Figure 4.
@@
 This might be a hint of an intrinsic correlation between dark matter-dark energy equality and the redshift where interaction energy transfer changes sign.
+Figure 5: Comparison of the effective equations of state in the interacting model and flat CPL.
+In Figure 5, we show that in the different models IntModel and fCPL, the effective equation of state of dark energy (left panel) is different but the total effective equation of state (right panel) is similar.
+This is because the observed data is more closely related to the total effective equation of state than the effective equation of state of dark energy.
+We now make a model comparison between the interacting model considered in this analysis (denoted IntModel), with the standard flat CPL parametrization (fCPL).
+Table 1: Comparison of model selection criteria. Each column lists the difference in the corresponding statistic relative to our interacting model, as in Equation 4.1.
+[…]
+  1.539   -4.934   3.536   8.918   -0.944
+In Table 1, we show the relative difference for a given statistic $S$, defined as
+$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=S(\text{IntModel})-S(\text{fCPL})\,.}$} (4.1)
+where [53, 37]
+$\Delta\ln\mathcal{Z}$ $=$ $\ln\mathcal{Z}_{\text{IntModel}}-\ln\mathcal{Z}_{\text{fCPL}}\,$, (4.2)
+$\mathrm{AIC}$ $=$ $-2\ln\mathcal{L}_{\mathrm{max}}+2k\,$, (4.3)
+$\mathrm{BIC}$ $=$ $-2\ln\mathcal{L}_{\mathrm{max}}+k\ln N\,$, (4.4)
+$\mathrm{DIC}$ $=$ $-2\ln\mathcal{L}(\bar{\theta})+2p_{D}\quad\mbox{where}~~p_{D}=-2~\overline{\ln\mathcal{L}(\theta)}+2\ln\mathcal{L}(\bar{\theta})\,.$ (4.5)
+The quantities above are defined as follows (see Appendix A for further details).
+  • $\mathcal{Z}$ is the Bayesian evidence, i.e., the marginal likelihood of the data given the model.
+  • $\mathcal{L}_{\mathrm{max}}$ is the maximum likelihood over the parameter space.
+  • $k$ is the total number of free parameters in the model.
+  • NN is the number of data points used in the fit.
+  • $\bar{\theta}$ is the posterior mean of the model parameters.
+  • $p_{D}$ is the effective number of parameters, quantifying model complexity.
+  • $\mathcal{L}(\theta)$ is the likelihood evaluated at a given parameter set $\theta$.
+  • $\overline{\ln\mathcal{L}(\theta)}$ is the posterior mean of the log-likelihood, which can be approximated as
+    $\overline{\ln\mathcal{L}(\theta)}\approx\frac{1}{N_{\mathrm{MC}}}\sum_{i=1}^{N_{\mathrm{MC}}}\ln\mathcal{L}(\theta_{i})\,$, (4.6)
+    where $\{\theta_{i}\}$ are the posterior samples and $N_{\mathrm{MC}}$ is the total number of samples.
+The preference of our IntModel compared to the fCPL model based on the different selection criteria is summarized as follows.
+  • fCPL provides a slightly better fit to the data according to $\chi^{2}$.
+  • Bayesian evidence shows moderate preference for the fCPL.
+  • AIC favours fCPL, though the difference is not statistically significant.
+  • BIC strongly prefers fCPL.
+  • DIC marginally favors IntModel, but the difference is negligible.
+Except for the BIC criterion, the IntModel has equivalent evidence to fCPL.
+The BIC favours fCPL because it strongly penalizes the IntModel, which has an extra parameter, $w_{q}$.
+In summary, the data does not show significant evidence for a preference of fCPL over IntModel.
 5 Conclusion
@@
 Then we reconstructed the redshift dependence of the interaction function, $Q(z)$.
-The parametrization of the intrinsic equation of state of dark energy behaves like thawing quintessence for $w_{\rm q}>-1$.
+The parametrization of the intrinsic equation of state of dark energy behaves like thawing quintessence for $w_{\rm q}>-1$, where $w_{\rm q}$ is the present value of the intrinsic equation of state of dark energy.
 On the other hand, it has phantom behavior for $w_{\rm q}<-1$.
@@
 […]
-In summary, we find interacting dark energy models can be viable alternatives to both the standard $\Lambda$CDM model and phantom dark energy scenarios, while the intrinsic dark energy equation of state is non-phantom.
+[…]
 Acknowledgments
 The authors are supported by the South African Radio Astronomy Observatory and the National Research Foundation (grant no. 75415).
+Appendix A Model selection criteria
+Here we briefly discuss the model selection criteria and decision rules associated with the statistical quantities $\chi^{2}$ and those defined in Equation 4.2Equation 4.5. We use the convention
+$\Delta S=S(\text{Model 1})-S(\text{Model 2})\,$, (A.1)
+so that positive (negative) values of $\Delta S$ mean the statistic is larger (smaller) for Model 1 than for Model 2.
+For each model comparison criterion, the tables below present the conventional scale for the strength of evidence for one model over another.
+The decision rules for different criteria are as follows.
+  • $\Delta\chi^{2}<0$, $\Delta\mathrm{AIC}<0$, $\Delta\mathrm{BIC}<0$, $\Delta\mathrm{DIC}<0$ all indicate stronger evidence in favour of Model 1 [43, 45, 31, 7, 49].
+  • $\Delta\ln\mathcal{Z}>0$ favours Model 1 [45, 31, 52].
+The quantitative criterion scale for model selection is listed in the tables below for each criterion.
+Table 2: Interpretation of $\Delta\ln\mathcal{Z}$ (Bayes evidence difference) according to the Jeffreys scale.
+  $\Delta\ln\mathcal{Z}$ Strength of evidence (Model 1 vs Model 2)
+  011 No substantial evidence for favouring Model 1 or 2
+  112.52.5 Weak evidence favouring Model 1 over 2
+  2.52.555 Moderate evidence favouring Model 1 over 2
+  >5>5 Strong evidence favouring Model 1 over 2
+Table 3: Interpretation of $\Delta\mathrm{AIC}=\mathrm{AIC}(\text{Model 1})-\mathrm{AIC}(\text{Model 2})$.
+  $\Delta\mathrm{AIC}$ Strength of evidence (Model 1 vs Model 2)
+  022 Model 1 slightly disfavoured relative to 2 (negligible to weak)
+  4477 Model 1 considerably disfavoured relative to 2 (moderate evidence against Model 1)
+  >10>10 Model 1 strongly disfavoured relative to 2 (essentially no support for Model 1)
+Table 4: Interpretation of $\Delta\mathrm{BIC}=\mathrm{BIC}(\text{Model 1})-\mathrm{BIC}(\text{Model 2})$.
+  $\Delta\mathrm{BIC}$ Strength of evidence (Model 1 vs Model 2)
+  022 Model 1 weakly disfavoured relative to 2
+  2266 Model 1 moderately disfavoured relative to 2
+  661010 Model 1 strongly disfavoured relative to 2
+  >10>10 Model 1 very strongly disfavoured relative to 2
+Table 5: Interpretation of $\Delta\mathrm{DIC}=\mathrm{DIC}(\text{Model 1})-\mathrm{DIC}(\text{Model 2})$.
+  $\Delta\mathrm{DIC}$ Strength of evidence (Model 1 vs Model 2)
+  022 Model 1 slightly disfavoured relative to 2
+  2277 Model 1 moderately disfavoured relative to 2
+  >10>10 Model 1 strongly disfavoured relative to 2

### v3

@@
 The effective equation of state of all the components of the model can be approximated as
-$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{w_{\rm tot}^{\rm eff}(z)=E^{-2}(z)\left[w^{\rm eff}_{\rm de}(z)\,\Omega_{{\rm de},0}\,f_{\rm de}(z)+w^{\rm eff}_{\rm dm}(z)\,\Omega_{{\rm dm},0}\,f_{\rm dm}(z)+\frac{1}{3}\Omega_{\rm r,0}(1+z)^{4}\right]\,,}$} (2.19)
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{w_{\rm tot}^{\rm eff}(z)=E^{-2}(z)\left[w^{\rm eff}_{\rm de}(z)\,\Omega_{{\rm de},0}\,f_{\rm de}(z)+w^{\rm eff}_{\rm dm}(z)\,\Omega_{{\rm dm},0}\,f_{\rm dm}(z)+\frac{1}{3}\Omega_{\rm r,0}(1+z)^{4}\right]\,,}$} (2.19)
 where $E(z)=H(z)/H_{0}$. At late times, we can neglect the radiation components and hence the last term in this equation.
@@
 In Table 1, we show the relative difference for a given statistic $S$, defined as
-$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=S(\text{IntModel})-S(\text{fCPL})\,.}$} (4.1)
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\Delta S=S(\text{IntModel})-S(\text{fCPL})\,.}$} (4.1)
 where [53, 37]